# Model Restrictiveness in Functional and Structural Settings

### v2

@@ Abstract
We generalize the notion of model restrictiveness in Fudenberg, Gao and Liang (2026) to a wider range of economic models with semi/non-parametric and structural ingredients. We show how restrictiveness can be defined and computed in infinite-dimensional settings using Gaussian process priors (including with shape restrictions) and other alternatives in Bayesian nonparametrics. We also extend the restrictiveness framework to structural models with endogeneity, instrumental variables, multiple equilibria, and nonparametric nuisance components. We discuss the importance of the user-specific choice of discrepancy functions in the context of Rademacher complexity and GMM criterion function, and relate restrictiveness to the limit of the average-case learning curve in machine learning. We consider applications to: (1) preferences under risk, (2) exogenous multinomial choice, and (3) multinomial choice with endogenous prices: for (1), we obtain results consistent with those in Fudenberg, Gao and Liang (2026); for (2) and (3), our findings show that nested logit and mixed logit exhibit similar restrictiveness under standard parametric specifications, and that IV exogeneity conditions substantially increase overall restrictiveness while altering model rankings.
Keywords: restrictiveness, complexity, flexibility, structural model, semiparametric, nonparametric, endogeneity, Bayesian nonparametric, Gausian process
@@ 1 Introduction
As Box (1976) famously remarked, “All models are wrong, but some are useful.” Almost all economic models are restrictive. They rule out some patterns in the data while emphasizing some others based on relevant economic theory, existing empirical knowledge, and intended usage of the model. The restrictiveness of economic models is a feature rather than a bug. Restrictions can encode economic structure, facilitate interpretation, and improve economic decision making. Yet restrictive models can be misspecified, and the nature of misspecification varies widely. Researchers often lack a quantitative sense of how much structure a given model imposes relative to plausible alternatives: for example, choosing between multinomial logit and mixed logit, or between different specifications of preferences under risk).
@@
Fudenberg, Gao and Liang (2026) develops a framework for measuring the restrictiveness of a (theoretical) model, viewed as a prediction rule $f:\mathcal{X}\to\mathcal{Y}$ mapping (exogenous) covariates $X$ to outcomes $Y$. Restrictiveness is defined relative to: (i) a user-specified set of eligible prediction rules $\mathcal{F}$, (ii) a discrepancy function $d$, and (iii) an evaluation distribution $\lambda_{\mathcal{F}}$ on $\mathcal{F}$. For a model class $\mathcal{F}_{\Theta}\subseteq\mathcal{F}$, restrictiveness measures the expected discrepancy between it and a pseudo-true rule drawn from $\lambda_{\mathcal{F}}$, normalized by the baseline discrepancy. Restrictiveness measures how much average approximation error is incurred by restricting to models in $\mathcal{F}_{\Theta}$ instead of using the flexible benchmark. Restrictiveness captures structural content but does not assess fit to observed data or statistical rejection-questions addressed by standard estimation, testing, and the completeness measure of Fudenberg, Kleinberg, Liang and Mullainathan (2022). Instead, restrictiveness quantifies how much structure a model imposes a priori.
@@
Restrictiveness has been adopted across several areas of economics as a tool for disciplined model evaluation. See, for example, Schwaninger (2022) on other-regarding preferences in dynamic bargaining, Ellis, Kariv and Ozbay (2024) on individual-level choice under uncertainty, Gentzkow, Shapiro, Yang and Yurukoglu (2024) on advertising pricing in media markets, and Ba, Bohren and Imas (2025) on over- and under- reaction in belief updating under cognitive constraints. Collectively, these applications illustrate the breadth of the restrictiveness–completeness framework: it provides a unified, quantitative language for comparing models across behavioral economics, industrial organization, and experimental game theory, enabling researchers to move beyond goodness-of-fit alone and toward a more principled assessment of the trade-off between empirical accuracy and theoretical discipline.
@@
A broader recent literature has also engaged with similar conceptual questions. de Clippel and Rozen (2024) addresses the closely related question of permissiveness, and emphasizes the concern that many bounded rationality theories can accommodate nearly arbitrary choice patterns. Andrews, Fudenberg, Lei, Liang and Wu (2025) studies a complementary dimension of model evaluation: transfer performance, or how well an economic model estimated in one domain predicts outcomes in another. Montiel Olea and Prat (2025) studies political competition in which candidates announce competing “ideologies”, and explicitly models the trade-off between data fit and model simplicity in an equilibrium setting.
Taken together, these diverse applications and conceptual engagements underscore the potential importance of the restrictiveness measure as a general-purpose tool in the evaluation of economic theories and models, which motivate us to further develop the restrictiveness measure beyond the finite-support and reduced-form setting of Fudenberg, Gao and Liang (2026).
@@
We then relate our notion of restrictiveness to existing measures of model complexity and discrepancy in statistics, econometrics, and machine learning. In particular, we examine the connection between restrictiveness and Rademacher complexity in a binary classification setting, and clarify that the degeneracy of restrictiveness established in Ellis and Neff (2025) is tied to a specific correlation-based discrepancy. We argue that the choice of discrepancy should be guided by interpretability and context rather than by existing capacity measures designed for other purposes. We also explain why GMM criterion functions are not suitable as discrepancy functions for restrictiveness: GMM criterion functions are defined as quadratic forms that capture violations of moment conditions, rather than measures of distance between the model-implied data distribution and the (pseudo-)true distribution. We further show that, in a pure approximation-error sense and in the absence of noise, restrictiveness can be interpreted as the normalized limit of the average-case learning curve, a well-studied concept in machine learning. A central message of our analysis is that $d$ is not fixed: it should be chosen to reflect the kind of approximation error that matters in the application, rather than inherited mechanically from existing complexity measures.
We apply our approach to three economic problems. First, we revisit the analysis of the restrictiveness of Cumulative Prospect Theory (CPT) and Disappointment Aversion (DA) models in Fudenberg, Gao and Liang (2026), replacing that paper’s finite set of binary lotteries with the entire space of monotone, bounded prediction rules for arbitrary binary lotteries. Second, we study the restrictiveness of standard discrete-choice models in industrial organization-multinomial logit, nested logit, and mixed logit, and quantify how much of mixed logit’s theoretical flexibility is actually realized by the parametric forms used in practice, and how the standard IO toolkit trades off restrictiveness and completeness. Third, we extend the analysis to multinomial choice with endogenous product characteristics. Using BLP-style instruments, we compare the discrete-choice models in an IV setting, where each specification is constrained both by its functional form and by moment conditions.
@@
Ellis and Neff (2025) studies the connection between restrictiveness and Rademacher complexity in a binary classification setting. It shows that, for a particular choice of eligible set and a discrepancy inspired by Rademacher complexity, a normalized version of our restrictiveness index is an affine transformation of the limiting Rademacher complexity of the model class. In that special case, all finite-dimensional falsifiable models appear “fully restrictive” in the limit. Section 4.1, interprets the degeneracy as a critique of this particular discrepancy rather than of restrictiveness itself: when $d$ is chosen to encode an existing capacity measure, restrictiveness inherits that measure’s asymptotic behavior and limitations. In addition, Ellis and Neff (2025) also propose a finite-sample version of discrepancy function $d_{n}$ (as a sample average using the $n$ data points on the features), which can be convenient in a variety of empirical settings. Ellis and Neff (2025) does not develop the asymptotic theory of $d_{n}$, which is required to construct confidence intervals for restrictiveness that reflects the finite-sample randomness of the covariates $X_{i}$; we provide that here.
Crawford (2025) uses Gaussian process priors to represent demand functions satisfying the restrictions of rational choice in consumer settings. It uses a binary (0-1) distance metric following Selten (1991) to evaluate whether observed demand can be rationalized exactly, as opposed to our use of a continuous discrepancy function. Continuous discrepancy functions help circumvent the degeneracy problem for conditional rationality that Crawford (2025) encounters when conditioning variables are continuously distributed. [^1] This degeneracy issue is related to the broader difficulty arising from 0-1 loss functions in VC dimension and Rademacher complexity analysis, as discussed in Ellis and Neff (2025).
@@
The rest of the paper is organized as follows. Section 2 describes how to define and compute restrictiveness in functional settings (with continuous feature space). Section 3 discusses how to define and compute the restrictiveness of structural econometric models with endogeneity, multiple equilibria, or semiparametric specifications. Section 4 relates and compares our restrictiveness measure to a variety of related but different existing concepts in statistics and econometrics. We consider three concrete applications in Section 5, and conclude in Section 6.
@@ 2.2 Evaluation and Numerical Implementation
Sampling GP with Monotonicity Constraints. We may want to sample functions that satisfy particular shape restrictions, e.g., boundedness, monotonicity, or convexity. Swiler, Gulian, Frankel, Safta and Jakeman (2020) surveys common strategies for incorporating constraints within Gaussian process regression. Our purpose is to sample from a constrained GP, which is equivalent to constrained GP regression without updating the prior. In general, there are two main categories - one enforces the constraints to hold globally through, for example, transforming the output of GP (Snelson, Ghahramani and Rasmussen, 2003) or imposing constraints on the coefficients of the spline functions (Maatouk and Bay, 2017; Shively et al., 2009); the other relaxes the global constraints to constraints at a finite set of “virtual” points (Riihimäki and Vehtari, 2010). In our later applications, the sampling algorithms we employ fall into the first category. We provide the details in Appendix F.
@@ 2.3 Estimation and Inference
Here, inference must account for both Monte Carlo uncertainty and sampling error in $d_{n}$. We derive the corresponding asymptotic distribution of $\widehat{r}_{n,M}$ for fixed $M$ as $n\to\infty$ and construct a feasible variance estimator that incorporates both sources of variability. The full procedure is provided in Appendix C.
@@ 3 Restrictiveness for Structural Models
So far we have treated $\mathcal{F}_{\Theta}$ as an abstract class of prediction rules. In structural economic models, however, the “structural function” mapping covariates to outcomes often involve strategic selections, interactions, or equilibrium that result in endogeneity and model incompleteness issues. In this section we explain how to define and compute restrictiveness for structural models.
@@ 3.1.1 Restrictiveness via Reduced-Form Representation
i.e., the restrictiveness of reduced-form model $\mathcal{F}_{\Theta,RF}$ under eligible set $\mathcal{F}$ based on the discrepancy function $d_{RF}$ according to Definition 1.
@@
researchers may only care about the mapping $f_{\theta}(x)$. In this case, we can simplify the definition of restrictiveness $r$. Specifically, we may take
  • $\mathcal{Y}$ to be the support of the outcome variable $Y_{i}$.
  • $\mathcal{F}$ to a given eligible set of mappings form $\mathcal{X}$ to $\mathcal{Y}$
  • $d$ to be the mean squared ($L_{2,X}$) distance.
We note that this simplification has also been considered in FGL, which proposes two “canonical” discrepancy functions, one based on KL divergence for distributions, and the other based on mean squared ($L_{2,X}$) distance for conditional expectation mappings. FGL also provides a discussion about why these two discrepancy functions are nicely “paired” with error functions underlying the definitions of completeness in Fudenberg, Kleinberg, Liang and Mullainathan (2022).
@@ 3.1.2 Restrictiveness under Structural-Form Error Additivity
Many econometric models are incomplete,in the sense that the structural form specificication may not admit a reduced-form representation as in Assumption 2. One class of such models are models with multiple equilibria, in which the reduced form takes the form of a correspondence rather than a function. We treat this scenario in Section 3.2, and show how the definition of restrictiveness can be adapted accordingly. Another important class of models without explicit reduced-form representations are partial equilibrium models identified via instrumental variables (IVs): for example, a demand model identified using exogenous demand and supply shifters without an explicit specification of the supply model. Such partial equilibrium models are prevalent in applied work, and often impose additivity of the structural error as in (9) below:
@@
for some known mappings $\Lambda$ and $f_{\theta}$ up to the parameter $\theta_{0}\in\Theta$, with $Y_{o,i}$ denoting the outcome variable and $Y_{c.i}$ denoting the endogenous covariates.
@@ 3.2 Structural Model with Multiple Equilibria
Definition in (12) implicitly assumes that the model $f$ has incorporated all relevant restrictions. Thus any $f\in\overline{\mathcal{F}}_{\Theta}$ is consistent with all stated model restrictions, which is why the definition uses the infimum approximation error. In some scenarios, one may want to impose an equilibrium selection restriction that picks a unique equilibrium. This additional equilibrium selection restriction effectively converts an incomplete model into a complete one and makes the minimization in equation (12) trivial: in other words, restrictiveness can be defined in the same way as in Definition 2.
@@
Consider the two-firm entry game as Tamer (2003):
@@ 3.3 Semiparametric Structural Models
Some structural models are semiparametric: some primitives are parametrically modeled, while others are left nonparametric. Often we may not wish to impose parametric assumptions on the unobserved error terms. A generic representation of such models takes the form of
@@
To numerically approximate $\mathcal{H}$ by simulating $M$ copies of $h$ from $\lambda_{\mathcal{H}}$, a specified space of functions mapping from $\mathcal{X}\times\mathcal{Z}$ to $\mathbb{R}^{J}$, and define

### v3

@@ Abstract
We extend the restrictiveness measure of Fudenberg, Gao and Liang (2026) to functional and structural econometric settings using Gaussian process priors. We find that models evaluated over continuum domains appear more restrictive than when evaluated over finite sets of observations. We also extend the restrictiveness framework to structural models with endogeneity, instrumental variables, multiple equilibria, and nonparametric nuisance components. We explain why the choice of discrepancy function is a substantive modeling decision, and why the Rademacher complexity and GMM criterion functions are unsuitable as discrepancies. We further show that restrictiveness equals the normalized limit of the noise-free average-case learning curve. In applications to preferences under risk, and multinomial choice under exogenous and endogenous settings, we find that the same models exhibit uniformly higher restrictiveness when evaluated over continuum domains than based on their predictions on finite sets, and that moment restrictions from endogeneity substantially increase restrictiveness and alter model rankings.
Keywords: restrictiveness, complexity, flexibility, structural model, semiparametric, nonparametric, endogeneity, Bayesian nonparametrics, Gaussian process.
@@ 1 Introduction
Economic models are restrictive by design: they rule out patterns in the data that conflict with theoretical priors, and this structure is what makes them useful for interpretation and decision making. But restrictive models can be misspecified, and researchers often lack a quantitative sense of how much structure a given model imposes relative to plausible alternatives. Should an applied IO economist use multinomial logit or mixed logit? How much additional structure does Cumulative Prospect Theory impose relative to Disappointment Aversion? These questions are important but difficult to answer without a formal measure of model restrictiveness. Because AI-assisted tools are making it easier to estimate vast numbers of candidate models on any given dataset (Yanagizawa-Drott, 2026), restrictiveness measures are more needed than ever.
@@
Fudenberg, Gao and Liang (2026) develops a framework for measuring the restrictiveness of a (theoretical) model, viewed as a prediction rule $f:\mathcal{X}\to\mathcal{Y}$ mapping (exogenous) covariates $X$ to outcomes $Y$. For a model class $\mathcal{F}_{\Theta}$ and a flexible benchmark $\mathcal{F}$, restrictiveness measures the expected approximation error incurred by restricting to $\mathcal{F}_{\Theta}$ instead of the benchmark, normalized by a baseline. Restrictiveness is complementary to goodness-of-fit: it captures the structural content of a model without reference to observed data, whereas fit-based measures such as the completeness index of Fudenberg et al. (2022) assess how well a model captures variation in an observed sample. Together, restrictiveness and completeness allow researchers to ask both how much a model rules out and how much of what matters it captures.
@@
Restrictiveness has been adopted across several areas of economics as a tool for disciplined model evaluation. See, for example, Schwaninger (2022) on other-regarding preferences in dynamic bargaining, Ellis, Kariv and Ozbay (2024) on individual-level choice under uncertainty, Gentzkow, Shapiro, Yang and Yurukoglu (2024) on advertising pricing in media markets, and Ba, Bohren and Imas (2025) on over- and under-reaction in belief updating under cognitive constraints. Collectively, these applications illustrate the breadth of the restrictiveness–completeness framework: it provides a unified, quantitative language for comparing models across behavioral economics, industrial organization, and experimental game theory, enabling researchers to make a principled assessment of the trade-off between empirical accuracy and theoretical discipline.
@@
A broader recent literature has engaged with related questions about model evaluation. de Clippel and Rozen (2024) addresses the closely related question of permissiveness, and emphasizes the concern that many bounded rationality theories can accommodate nearly arbitrary choice patterns. Andrews, Fudenberg, Lei, Liang and Wu (2025) studies a complementary dimension of model evaluation: transfer performance, or how well an economic model estimated in one domain predicts outcomes in another. Montiel Olea and Prat (2025) studies political competition in which candidates announce competing “ideologies”, and explicitly models the trade-off between data fit and model simplicity in an equilibrium setting.
@@
We show that the Rademacher complexity and VC dimension measures encode a specific discrepancy function that is poorly suited for measuring economic model restrictiveness, and relate this to the degeneracy result of Ellis and Neff (2025). We argue that the choice of discrepancy should be guided by interpretability and context rather than by existing capacity measures designed for other purposes. We also explain why GMM criterion functions are not suitable as discrepancy functions for restrictiveness: GMM criterion functions are defined as quadratic forms that capture violations of moment conditions, rather than measures of distance between the model-implied data distribution and the (pseudo-)true distribution. We further show that, in a pure approximation-error sense and in the absence of noise, restrictiveness can be interpreted as the normalized limit of the average-case learning curve, a well-studied concept in machine learning. A central message of our analysis is that $d$ is not fixed: it should be chosen to reflect the kind of approximation error that matters in the application, rather than inherited mechanically from existing complexity measures.
We apply our approach to three economic problems. First, we revisit the analysis of the restrictiveness of Cumulative Prospect Theory (CPT) and Disappointment Aversion (DA) models in Fudenberg, Gao and Liang (2026), replacing that paper’s finite set of binary lotteries with the entire space of monotone, bounded prediction rules for arbitrary binary lotteries. Second, we study the restrictiveness of standard discrete-choice models in industrial organization (multinomial logit, nested logit, and mixed logit), and quantify how much of mixed logit’s theoretical flexibility is actually realized by the parametric forms used in practice, and how the standard IO toolkit trades off restrictiveness and completeness. Third, we extend the analysis to multinomial choice with endogenous product characteristics. Using BLP-style instruments, we compare the discrete-choice models in an IV setting, where each specification is constrained both by its functional form and by moment conditions.
Computing restrictiveness is practically feasible: whenever a replication package exists to reproduce empirical results, the same computational pipeline can compute restrictiveness, requiring only the specification of an elible set $\mathcal{F}$, a discrepancy function $d$, and a Bayesian nonparametric prior as the evaluation distribution $\lambda_{\mathcal{F}}$. As we argue in this paper, these are precisely the objects that researchers should contemplate depending on their specific context.
@@
Ellis and Neff (2025) studies the connection between restrictiveness and Rademacher complexity in a binary classification setting. It shows that, for a particular choice of eligible set and a discrepancy inspired by Rademacher complexity, a normalized version of our restrictiveness index is an affine transformation of the limiting Rademacher complexity of the model class. In that special case, all finite-dimensional falsifiable models appear “fully restrictive” in the limit. Section 4.1 interprets the degeneracy as a critique of this particular discrepancy rather than of restrictiveness itself: when $d$ is chosen to encode an existing capacity measure, restrictiveness inherits that measure’s asymptotic behavior and limitations. In addition, Ellis and Neff (2025) also propose a finite-sample version of discrepancy function $d_{n}$ (as a sample average using the $n$ data points on the features), which can be convenient in a variety of empirical settings. Ellis and Neff (2025) does not develop the asymptotic theory of $d_{n}$, which is required to construct confidence intervals for restrictiveness that reflect the finite-sample randomness of the covariates $X_{i}$; we provide that here.
Crawford (2025) uses Gaussian process priors to represent demand functions satisfying the restrictions of rational choice in consumer settings. It uses a binary (0-1) distance metric following Selten (1991) to evaluate whether observed demand can be rationalized exactly, as opposed to our use of a continuous discrepancy function. Continuous discrepancy functions help circumvent the degeneracy problem for conditional rationality that Crawford (2025) encounters when conditioning variables are continuously distributed. [^1] This degeneracy issue is related to the broader difficulty arising from 0-1 loss functions in VC dimension and Rademacher complexity analysis, as discussed in Ellis and Neff (2025). Hansen and Jagannathan (1997) measure misspecification with least-squares distance, as in our discussion of GMM estimation in Section 4.2. As we explain there, however, this criterion depends on the available instruments and the observed data in ways that restrictiveness does not.
@@
Section 2 describes how to define and compute restrictiveness in functional settings (with continuous feature space). Section 3 discusses how to define and compute the restrictiveness of structural econometric models with endogeneity, multiple equilibria, or semiparametric specifications. Section 4 relates and compares our restrictiveness measure to a variety of related but different existing concepts in statistics and econometrics. We consider three concrete applications in Section 5, and conclude in Section 6.
@@ 2.2 Evaluation and Numerical Implementation
Sampling GP with Monotonicity Constraints. We may want to sample functions that satisfy particular shape restrictions, e.g., boundedness, monotonicity, or convexity. Swiler, Gulian, Frankel, Safta and Jakeman (2020) surveys common strategies for incorporating constraints within Gaussian process regression. Our purpose is to sample from a constrained GP, which is equivalent to constrained GP regression without updating the prior. In general, there are two main categories - one enforces the constraints to hold globally through, for example, transforming the output of GP (Snelson, Ghahramani and Rasmussen, 2003) or imposing constraints on the coefficients of the spline functions (Maatouk and Bay, 2017; Shively et al., 2009); the other relaxes the global constraints to constraints at a finite set of “virtual” points (Riihimäki and Vehtari, 2010). In our later applications, the sampling algorithms we employ fall into the first category. We provide the details in Online Appendix S.4.
@@ 2.3 Estimation and Inference
Here, inference must account for both Monte Carlo uncertainty and sampling error in $d_{n}$. We derive the corresponding asymptotic distribution of $\widehat{r}_{n,M}$ for fixed $M$ as $n\to\infty$ and construct a feasible variance estimator that incorporates both sources of variability. The full procedure is provided in Appendix B.
@@ 3 Restrictiveness for Structural Models
So far we have treated $\mathcal{F}_{\Theta}$ as an abstract class of prediction rules. This section extends restrictiveness to structural models in three scenarios, each requiring different treatment: (1) models with explicit reduced forms, (2) partial-equilibrium models identified via instruments, and (3) models with incomplete structure, e.g.multiple equilibria or semiparametric components.
@@ 3.1.1 Restrictiveness via Reduced-Form Representation
i.e., the restrictiveness of reduced-form model $\mathcal{F}_{\Theta,RF}$ under eligible set $\mathcal{F}_{RF}$ based on the discrepancy function $d_{RF}$ according to Definition 1.
@@
researchers may only care about the mapping $f_{\theta}(x)$. In this case, the definition of restrictiveness admits a natural simplification. We take $\mathcal{Y}$ to be the support of the $Y_{i}$, $\mathcal{F}$ to be an eligible set of mappings from $\mathcal{X}$ to $\mathcal{Y}$, and $d$ to be the mean squared $L_{2,X}$ As noted in FGL Appendix E, this discrepancy function pairs naturally with mean squared error, which is typically used in the definition of completeness.
@@ 3.1.2 Restrictiveness under Structural-Form Error Additivity
Many econometric models are incomplete, in the sense that the structural form specification may not admit a reduced-form representation as in Assumption 2. One class of such models are models with multiple equilibria, in which the reduced form takes the form of a correspondence rather than a function. We treat this scenario in Section 3.2, and show how the definition of restrictiveness can be adapted accordingly. Another important class of models without explicit reduced-form representations are partial equilibrium models identified via instrumental variables (IVs): for example, a demand model identified using exogenous demand and supply shifters without an explicit specification of the supply model. Such partial equilibrium models are prevalent in applied work, and often impose additivity of the structural error as in (9) below:
@@
for some known mappings $\Lambda$ and $f_{\theta}$ up to the parameter $\theta_{0}\in\Theta$, with $Y_{o,i}$ denoting the outcome variable and $Y_{c,i}$ denoting the endogenous covariates.
@@ 3.2 Structural Model with Multiple Equilibria
The definition in (12) implicitly assumes that the model $f$ has incorporated all relevant restrictions. Thus any $f\in\overline{\mathcal{F}}_{\Theta}$ is consistent with all stated model restrictions, which is why the definition uses the infimum approximation error. In some scenarios, one may want to impose an equilibrium selection restriction that picks a unique equilibrium. This additional equilibrium selection restriction effectively converts an incomplete model into a complete one and makes the minimization in equation (12) trivial: in other words, restrictiveness can be defined in the same way as in Definition 2.
@@
Consider the two-firm entry game in Tamer (2003):
@@ 3.3 Semiparametric Structural Models
Some structural models are semiparametric: some primitives are parametrically modeled, while others are left nonparametric. Often we may not wish to impose parametric assumptions on the unobserved error terms. A generic representation of such models takes the form
@@
We numerically approximate $\mathcal{H}$ by simulating $M$ copies of $h$ from $\lambda_{\mathcal{H}}$, a specified space of functions mapping from $\mathcal{X}\times\mathcal{Z}$ to $\mathbb{R}^{J}$, and define